Jeremy sells a magazine which is produced in order to raise money for the homeless people. The probability of making a sale is independently ,0.09 for each person he approaches. Given that he approaches 40 people, find the probability that he will make

I) exactly 4 sales

ii) at most 3 sales

iii)more than 3 sales



1
Expert's answer
2021-08-10T11:13:50-0400

"p=0.09 \\

q=1-p=1-0.09=0.91 \\

n=40 \\

P(X=k) = C^n_k \times p^k \times q^{n-k}"

i)

"P(X=4) = \frac{40!}{4!(40-4)!} \times 0.09^4 \times 0.91^{40-4} \\

= 91390 \times 6.561 \times 10^{-5} \times 0.033534 \\

= 0.2010"

ii)

"P(X≤3) = P(X=0) + P(X=1) + P(X=2) +P(X=3) \\

P(X=0) = \frac{40!}{0!(40-0)!} \times 0.09^0 \times 0.91^{40-0} \\

= 1 \times 1 \times 0.02299 \\

= 0.02299 \\

P(X=1) = \frac{40!}{1!(40-1)!} \times 0.09^1 \times 0.91^{40-1} \\

= 40 \times 0.09 \times 0.0252705 \\

= 0.09097 \\

P(X=2) = \frac{40!}{2!(40-2)!} \times 0.09^2 \times 0.91^{40-2} \\

= 780 \times 0.0081 \times 0.027769 \\

= 0.17544 \\

P(X=3) = \frac{40!}{3!(40-3)!} \times 0.09^3 \times 0.91^{40-3} \\

= 9880 \times 0.000729 \times 0.030516 \\

= 0.21979 \\

P(X≤3) = 0.02299+ 0.09097+ 0.17544 +0.21979=0.50919"

iii)

"P(X>3) = 1 -P(X≤3) \\

= 1 -0.50919 \\

= 0.49081"